# RuleArena: A Benchmark for Rule-Guided Reasoning with LLMs in Real-World Scenarios

### v1

@@ Abstract
This paper introduces RuleArena, a novel and challenging benchmark designed to evaluate the ability of large language models (LLMs) to follow complex, real-world rules in reasoning. Covering three practical domains—airline baggage fees, NBA transactions, and tax regulations—RuleArena assesses LLMs’ proficiency in handling intricate natural language instructions that demand long-context understanding, logical reasoning, and accurate mathematical computation. Two key attributes distinguish RuleArena from traditional rule-based reasoning benchmarks: (1) it extends beyond standard first-order logic representations, and (2) it is grounded in authentic, practical scenarios, providing insights into the suitability and reliability of LLMs for real-world applications. Our findings reveal several notable limitations in LLMs: (1) they struggle to identify and apply the appropriate rules, frequently becoming confused by similar but distinct regulations, (2) they cannot consistently perform accurate mathematical computations, even when they correctly identify the relevant rules, and (3) in general, they perform poorly in the benchmark. These results highlight significant challenges in advancing LLMs’ rule-guided reasoning capabilities in real-life applications.
@@ 1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Team, 2023; Anthropic, 2024) have demonstrated remarkable capabilities across a range of real-world applications, including code synthesis (Rozière et al., 2023) and customer service (Shi et al., 2024). However, their limited domain-specific knowledge often leads to generating unfaithful or misleading information, which can cause significant risks and financial liabilities. For example, Canadian airline was recently required to compensate a customer who received incorrect guidance from the airline’s chatbot [^1] https://www.theguardian.com/world/2024/feb/16/air-canada-chatbot-lawsuit. These challenges highlight the need for robust, real-world benchmarks that assess how faithfully and accurately LLMs can follow real-life instructions and adhere to relevant regulations, thereby ensuring reliable and safe outputs for deployment.
Although several studies have examined LLMs’ instruction-following abilities (Chen et al., 2024a; Jiang et al., 2024; Wen et al., 2024), they have primarily focused on stylistic constraints, such as the expected format Zhou et al. (2023), length, or topic of responses. Yet, the significance of instruction-following extends well beyond stylistic compliance. In many problem-solving scenarios, instructions function as rules: they impose logical constraints on the reasoning process and specify how answers should be derived from given inputs. However, limited attention has been devoted to LLMs’ capacity to follow complex rules.
@@
As illustrated in Figure 1, RuleArena is developed from three representative, real-world scenarios: (1) airline luggage payment regulations, (2) NBA transactions, including player acquisition and player transfers, and (3) taxation policies. From these domains, we collect authentic rules currently implemented by companies or government agencies. For each domain, we construct a set of challenging test problems, pairing each question with a ground-truth solution, and then evaluate a range of state-of-the-art LLMs on their ability to conform to the rules. Provided with the domain-specific task instructions and reference rules, LLMs are required to resolve each test problem through reasoning according to the information from the question under the guidance of reference rules.
@@
  • A comprehensive analysis of prevalent challenges: By examining common failure cases and identifying difficult rule types, we uncover several systematic issues that limit current LLMs’ rule-guided reasoning capabilities.
Upon extensive experiments on RuleArena, our main findings are as follows:
  • Existing state-of-the-art LLMs, such as GPT-4o and Claude-3.5 Sonnet, mostly fail on complex rule-guided reasoning tasks.
  • LLMs struggle integrate multiple rules or facts cohesively and are prone to distraction by irrelevant information.
  • Further analyses identify common failure modes, including inadequate rule recall, improper usage of similar rules, and computational errors.
@@ 2 Related Work
A wide range of benchmarks has been designed to evaluate LLMs’ instruction-following abilities from various perspectives, including semantics (Zheng et al., 2023; Li et al., 2023; Liu et al., 2023), format (Xia et al., 2024; Tang et al., 2023), and response length (Chen et al., 2024b; Sun et al., 2023). To further probe complexity, some works have introduced benchmarks that construct complex instructions through compositional methods. For example, WizardLM (Xu et al., 2023) generates intricate tasks by combining simpler instructions, while CELLO (He et al., 2024) uses task descriptions and input texts to create complex prompts grounded in real-world scenarios. ComplexBench (Wen et al., 2024) adopts multiple compositional structures to integrate atomic requirements into more challenging instructions. In contrast, our work focuses on instructions derived directly from real-life scenarios, where naturally occurring complexities arise from multifaceted constraints incurred by inputs on the set of instructions.
@@ 3.1 Domains and Rule Collection
We select three real-life domains that are both familiar in everyday life and demonstrate a high level of complexity:
@@ 3.3 Difficulty control
To assess LLMs’ capabilities under varying levels of complexity, we create problems with different degrees of difficulty. We define three levels of difficulties in each domain:
@@ 3.4 Evaluation Metrics
The first category focuses on problem-level evaluations: for each problem, we examine whether all necessary rules were applied, whether any extraneous rules were applied, and whether the final answer aligns with the ground-truth solution:
@@
The second category of metrics focuses on rules rather than problems. For each rule in the domain, we assess whether it is applied to all problems that require it, and whether the problems it is applied to are exactly those that necessitate its use:
@@
Table 2: Main problem-wise evaluation results on airline, NBA, and tax domains. ${\rm P}(t)$ denotes problem-wise precision, ${\rm AC}(t)$ denotes problem-wise rule application correctness, and ${\rm R}(t)$ denotes problem-wise recall.
@@ 4.1 Experiment Settings
LLMs. Our rules, which are prompted directly into LLMs, can be of a length up to 20,000 tokens. Therefore, we only consider LLMs that can handle such long contexts, including Llama-3.1 70B, Llama-3.1 405B (Dubey et al., 2024), Qwen-2.5 72B (Qwen Team, 2024), Claude-3.5 Sonnet (Anthropic, 2024), and GPT-4o (OpenAI, 2024)  [^6] Currently we have limited access to o1 API and thus only partial result, and we will update o1 performance once we have full access to the API.
@@
Output Parsing. To obtain the rule-usage matrix $U$ we mentioned in Section 3.4, we utilize the structured output mode of GPT-4o (OpenAI, 2024) to parse the raw textual responses from LLMs. Specifically, for airline and tax problems we structuralize the ground-truth calculation process and ask GPT-4o to fill in problem-specific information according to an LLM’s response, while for NBA problems we enumerate a list of all rules and ask GPT-4o to directly judge whether a specific rule is applied in an LLM’s response. For details we refer our readers to Appendix C.
@@ 4.2.1 Problem-wise Analysis
Low Accuracy. Overall performance in problem result accuracy (${\rm Acc}(t)$), as summarized in Table 2, remains unsatisfactory across all three scenarios. Under the 0-shot setting, even advanced models such as Llama 405B, Claude-3.5, and GPT-4o fail to produce correct answers for the simplest test problems. For more challenging problems, particularly in the airline and tax domains, ${\rm Acc}(t)$ rarely exceeds 10%. In 1-shot setting, we notice marked improvements on the easiest problems, yet the gains diminish as problem difficulty increases. These persistently low ${\rm Acc}(t)$ highlight the inherent complexity of the RuleArena benchmark and emphasize the need for more robust LLM reasoning and rule-following capabilities.
@@
High Rule Application Correctness. While LLMs demonstrate relatively application correctness (${\rm AC}(t)$) on rule application computation on average, ${\rm AC}(t)$ never reaches a perfect 100%. Occasional errors in mathematical calculations or logical operations emerge even under explicit rule guidance. Although these mistakes are not pervasive, a single computational error can significantly compromise the final output’s accuracy (${\rm Acc}(t)$) in many cases. This observation underscores the importance of improving the reliability of math computation abilities in LLMs.
@@ 4.2.2 Rule-wise Analysis
Here, we provide a detailed examination of the rule-level evaluation. Figure 2 presents recall (${\rm R}(r)$) and application correctness (${\rm AC}(r)$) in airline domain and metrics for NBA transaction and tax domains are presented in Figure 7 and Figure 8 in Appendix. Table 3 summarizes the metric results by reporting the mean and variance of three key metrics across all rules: recall (${\rm R}(r)$), application correctness (${\rm AC}(r)$), and precision (${\rm P}(r)$). The low variance observed in metrics such as ${\rm P}(r)$ within the airline and tax domains suggests that certain performance aspects are largely independent of the specific rules being applied. In contrast, the high variance seen in metrics like ${\rm R}(r)$ implies that recall performance is significantly influenced by the particular rules in question. Detailed analysis is presented below.
@@
(a) Airline
(b) NBA
(c) Tax
5 What Impacts Rule Following?
5.1 Correlation Between Accuracy and Other Metrics
5.2 Do In-Context Examples Help?
Rule Setting ${\rm R}(r)$ ${\rm P}(r)$
Over 38 rule 0-shot 0.00 N/A
1-shot 0.35 0.76
[…]
1-shot 0.23 0.20
Table 7: 0-shot and 1-shot rule-wise comparison.
5.3 Does Rule Representation Matter?
Models Setting Airline Tax
${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$ ${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$
Llama 70 Table 0.764 0.558 0.01 0.834 0.989 0.01
Text 0.764 0.582 0.01 0.814 0.991 0.00
[…]
Text 0.748 0.633 0.02 0.859 0.996 0.01
[…]
Text 0.835 0.587 0.07 0.919 0.998 0.05
[…]
Text 0.937 0.705 0.06 0.971 1.000 0.33
[…]
Text 0.864 0.669 0.03 0.960 1.000 0.33
(a) Correctness
(b) Recall
(c) Accuracy
5.4 Do Distractive Rules Matter?
5.5 Intermediate Summary
In summary, various factors—such as rule complexity, the presence of distractive information, and the difficulty gap between in-context examples and target problems—can profoundly influence LLM performance. Even when LLMs succeed in simpler conditions, challenges like complex mathematical reasoning, large amounts of extraneous rules, and non-ideal in-context samples can severely limit their effectiveness on RuleArena problems.
6 Case Studies
@@ 7 Conclusions
In this paper, we introduce RuleArena, a real-world benchmark designed to evaluate the abilities of LLMs on various rule-guided reasoning tasks. We observe that existing LLMs face significant challenges when they try to tackle problems on RuleArena - even the strongest Claude-3.5 and GPT-4o models can hardly succeed on our hardest tasks. Our further analysis indicates that LLMs struggle to integrate multiple rules or facts cohesively and are prone to irrelevant distractions.
RuleArena’s intricate design and emphasis on real-world rule-based reasoning represent an essential step in advancing LLMs’ capabilities. By posing fundamental challenges in complex rule recall and multi-rule integration, this benchmark provides a valuable tool for understanding and enhancing the reasoning abilities of LLMs. We envision RuleArena as a foundation for future research that strives to improve LLM performances in solving increasingly complex tasks.
8 Future Directions
While this study provides a comprehensive evaluation and analysis of LLMs’ rule-guided reasoning capabilities, there remain numerous promising avenues for future research:
@@ 8 Future Directions
In this work, we rely on GPT-4o to parse textual responses into structured JSON, facilitating downstream analyses of rule application. A logical next step would be to investigate the use of LLMs for fully automated reasoning evaluations, including the identification of intermediate errors. This direction aligns with the concept of “LLM-as-a-judge” (Zheng et al., 2023), which, despite potential bias or inaccuracies (Xu et al., 2024), offers a scalable alternative to labor-intensive human evaluation and could improve the reliability and granularity of assessment metrics.
@@
Supervised fine-tuning has proven effective in enhancing LLMs’ performance on tasks requiring substantial domain knowledge (Jeong, 2024; Fu et al., 2023; Wu et al., 2024). While we did not pursue fine-tuning in this study—given the high cost of obtaining extensive rule-guided reasoning data and the limited generalizability to unseen domains—it remains a worthwhile direction. Investigating whether training with datasets from related domains, such as mathematical or logical reasoning tasks or code generation problems, can bolster LLMs’ rule-following performance is an open question worth exploring.
Improving Rule Recall and Aggregation
Our experiments reveal that LLMs frequently struggle with recalling and aggregating the correct rules, and that problem-wise recall strongly correlates with overall accuracy. Addressing these challenges may involve refining rule retrieval mechanisms or integrating structured reasoning frameworks. For instance, approaches that convert rules into structured data (Pan et al., 2023; Wang et al., 2024) have shown promise in reasoning tasks. Building on these insights, a hybrid system that combines LLM-based reasoning with symbolic reasoners may enhance both the consistency and robustness of rule-guided reasoning in real-world scenarios.

### v2

@@ Abstract
This paper introduces RuleArena, a novel and challenging benchmark designed to evaluate the ability of large language models (LLMs) to follow complex, real-world rules in reasoning. Covering three practical domains—airline baggage fees, NBA transactions, and tax regulations—RuleArena assesses LLMs’ proficiency in handling intricate natural language instructions that demand long-context understanding, logical reasoning, and accurate mathematical computation. Two key attributes distinguish RuleArena from traditional rule-based reasoning benchmarks: (1) it extends beyond standard first-order logic representations, and (2) it is grounded in authentic, practical scenarios, providing insights into the suitability and reliability of LLMs for real-world applications. Our findings reveal several notable limitations in LLMs: (1) they struggle to identify and apply the appropriate rules, frequently becoming confused by similar but distinct regulations, (2) they cannot consistently perform accurate mathematical computations, even when they correctly identify the relevant rules, and (3) in general, they perform poorly in the benchmark. We also observe a significant performance boost when LLMs are provided with external tools for oracle math and logic operations. These results highlight significant challenges and promising research directions in advancing LLMs’ rule-guided reasoning capabilities in real-life applications. Our codes and data are publicly available on GitHub.
@@ 1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Team, 2023; Anthropic, 2024) have demonstrated remarkable capabilities across many real-world applications. However, their limited domain-specific knowledge often leads to unfaithful or misleading output, which can cause significant risks and financial liabilities. For example, Canadian airline was recently required to compensate a customer who received incorrect guidance from the airline’s chatbot [^1] https://www.theguardian.com/world/2024/feb/16/air-canada-chatbot-lawsuit. These challenges highlight the need for robust, real-world benchmarks that assess how faithfully and accurately LLMs can follow real-life instructions and adhere to relevant regulations, thereby ensuring reliable and safe outputs for deployment.
Although several studies have examined LLMs’ instruction-following abilities (Chen et al., 2024a; Jiang et al., 2024; Wen et al., 2024), they mainly focused on style constraints, such as the format Zhou et al. (2023), length, or topic of responses. Yet, the significance of instruction-following extends well beyond style compliance. In many problem-solving scenarios, instructions function as rules: they impose logical constraints on the reasoning process and specify how answers should be derived from given inputs. However, limited attention has been paid to LLMs’ capacity to follow complex rules.
@@
As illustrated in Figure 1, RuleArena is developed from three representative, real-world scenarios: (1) airline luggage fee policies, (2) NBA transactions, and (3) taxation policies. From these domains, we collect authentic rules currently implemented by companies or government agencies. For each domain, we construct a set of challenging test problems, pairing each question with a ground-truth solution, and then evaluate a range of state-of-the-art LLMs on their ability to conform to the rules. LLMs are provided with the domain-specific task instructions and reference rules, and required to resolve each test problem through reasoning according to the question and reference rules.
@@
  • A comprehensive analysis of prevalent challenges: By examining common failure cases, identifying difficult rule types, and conducting extensive controlled experiments, we uncover several systematic issues that limit current LLMs’ rule-guided reasoning capabilities and promising directions to improve.
@@ 2 Related Work
A wide range of benchmarks has been designed to evaluate LLMs’ instruction-following abilities from various perspectives, including semantics (Zheng et al., 2023; Li et al., 2023; Liu et al., 2023; Wu et al., 2024a, b), format (Xia et al., 2024; Tang et al., 2023), and response length (Chen et al., 2024b; Sun et al., 2023). To further probe complexity, some works have introduced benchmarks that construct complex instructions through compositional methods. For example, WizardLM (Xu et al., 2023) generates intricate tasks by combining simpler instructions, while CELLO (He et al., 2024) uses task descriptions and input texts to create complex prompts grounded in real-world scenarios. ComplexBench (Wen et al., 2024) adopts multiple compositional structures to integrate atomic requirements into more challenging instructions. In contrast, our work focuses on instructions derived directly from real-life scenarios, where naturally occurring complexities arise from multifaceted constraints incurred by inputs on the set of instructions.
@@ 3.1 Domains and Rule Collection
We select three real-life domains that all both familiar in everyday life and demonstrate a high level of complexity:
@@ 3.3 Difficulty control
To assess LLMs’ capabilities under varying levels of complexity, we create problems with different degrees of difficulty. We define three levels of difficulties in each domain.
@@
The statistics of problems at different difficulty levels in each domain are listed in Table 2.
Airline NBA Tax
Level 1 100 81 100
Level 2 100 89 100
Level 3 100 46 100
In Total 300 216 300
Table 2: Number of test problems at different difficulty levels in each domain.
@@ 3.4 Evaluation Metrics
The first group focuses on problem-level evaluations: for each problem, we examine whether all necessary rules were applied, whether any extraneous rules were applied, and whether the final answer aligns with the ground-truth solution:
@@
The second group of metrics focuses on rules rather than problems. For each rule in the domain, we assess whether it is applied to all problems that require it, and whether the problems it is applied to are exactly those that necessitate it:
@@
[…]
1-shot 1.000 0.994 1.000 0.68 1.000 0.960 0.994 0.33 1.000 0.894 0.939 0.24
Table 3: Main problem-wise evaluation results on airline, NBA, and tax domains. ${\rm P}(t)$ denotes problem-wise precision, ${\rm AC}(t)$ denotes problem-wise rule application correctness, and ${\rm R}(t)$ denotes problem-wise recall. The best and second best results in each column are rendered with orange and blue backgrounds respectively.
@@ 4.1 Experiment Settings
LLMs. Our rules, which are prompted directly into LLMs, can be of a length up to 20,000 tokens. Therefore, we only consider LLMs that can handle such long contexts, including Llama-3.1 70B, Llama-3.1 405B (Dubey et al., 2024), Qwen-2.5 72B (Qwen Team, 2024), Claude-3.5 Sonnet (Anthropic, 2024), GPT-4o (OpenAI, 2024a), and o1-preview (OpenAI, 2024b).
@@
Output Parsing. To obtain the rule-usage matrix $U$ we mentioned in Section 3.4, we utilize the structured output mode of GPT-4o (OpenAI, 2024a) to parse the raw textual responses from LLMs. Specifically, for airline and tax problems we structuralize the ground-truth calculation process and ask GPT-4o to fill in problem-specific information according to an LLM’s response, while for NBA problems we enumerate a list of all rules and ask GPT-4o to directly judge whether a specific rule is applied in an LLM’s response. For details we refer our readers to Appendix C.
@@ 4.2.1 Problem-wise Analysis
Low Accuracy. Overall performance in problem result accuracy (${\rm Acc}(t)$), as summarized in Table 3, remains unsatisfactory across all three scenarios. Under the 0-shot setting, non-reasoning LLMs such as Llama 405B, Claude-3.5, and GPT-4o fail to produce correct answers for the simplest test problems, and even advanced reasoning model like o1-preview can solve only about 50$\sim$60of Level 1 problems. For more challenging problems, particularly in the airline and tax domains, ${\rm Acc}(t)$ of non-reasoning models rarely exceeds 10% and o1-preview fails most of them as well. In 1-shot setting, we notice marked improvements on the easiest problems, yet the gains diminish as problem difficulty increases. These persistently low ${\rm Acc}(t)$ highlight the inherent complexity of the RuleArena benchmark and emphasize the need for more robust LLM reasoning and rule-following capabilities.
@@
High Rule Application Correctness. While LLMs demonstrate relatively high application correctness (${\rm AC}(t)$) on rule application computation on average, ${\rm AC}(t)$ never reaches a perfect 100%. Occasional errors in mathematical calculations or logical operations emerge even under explicit rule guidance. Although these mistakes are not pervasive, a single computational error can significantly compromise the final output’s accuracy (${\rm Acc}(t)$) in many cases. This observation underscores the importance of improving the reliability of math computation abilities in LLMs.
@@ 4.2.2 Rule-wise Analysis
Here, we provide a detailed examination of the rule-level evaluation. Figure 2 presents recall (${\rm R}(r)$) and application correctness (${\rm AC}(r)$) in airline domain and metrics for NBA transaction and tax domains are presented in Figure 5 and Figure 6 in Appendix. Table 4 summarizes the metric results by reporting the mean and variance of three key metrics across all rules: recall (${\rm R}(r)$), application correctness (${\rm AC}(r)$), and precision (${\rm P}(r)$). The low variance observed in metrics such as ${\rm P}(r)$ within the airline and tax domains suggests that certain performance aspects are largely independent of the specific rules being applied. In contrast, the high variance seen in metrics like ${\rm R}(r)$ implies that recall performance is significantly influenced by the particular rules in question. Detailed analysis is presented below in Table 4.
@@
4.3 In-Depth Analyses
@@
4.3.1 What Impacts Rule Following?
We study the factors influencing LLM performance, as measured by ${\rm Acc}(t)$, and provide the complete experiment results in Appendix D.2. The main findings are:
Correlation between ${\rm Acc}(t)$ and other metrics. We compare the correlation between problem-wise metrics (i.e., ${\rm P}(t)$, ${\rm AC}(t)$, ${\rm R}(t)$) and accuracy ${\rm Acc}(t)$. The correlation is the most obvious and almost linear between ${\rm R}(t)$ and ${\rm Acc}(t)$, while highly non-linear or unclear between other two metrics and ${\rm Acc}(t)$.
The effect of in-context examples. We observe that LLMs generally provide better performances given 1-shot example on airline, tax, and (easy) NBA problems. However, when tackling more challenging NBA problems (Levels 2 and 3), providing an example increases ${\rm P}(t)$ and ${\rm R}(t)$ but leads to a counterintuitive decrease in overall ${\rm Acc}(t)$.
Rule representation has a mild effect. In the airline and tax domains, some rules are represented as Markdown tables. To test whether representation format affects performance, we convert these tabular rules into textual “if-then” statements and compare with original results. The comparison shows that converting tabular rules into text improves ${\rm R}(r)$, but has little impact on other metrics, including ${\rm Acc}(t)$.
Distractive rules degrades LLM performance. An essential aspect of rule-following involves identifying which rules are relevant to the current problem. We assess the extent to which irrelevant rules detrimentally affect performance, and notice that the presence of distractive (irrelevant) rules significantly degrades LLM performance.
Tool augmentation boosts overall performance. Using external tools is a simple way to reduce math and logic errors from LLMs. To study to what extent external tools can help LLM rule-guided reasoning, we ask our LLMs to write Python code and use the execution result as solution, where Python interpreter serves as an oracle math and logic calculator. We observe that LLMs can achieve a significant performance boost with tool augmentation but are still far from perfect.
4.3.2 Case Studies
@@ 5 Conclusions
In this paper, we introduce RuleArena, a real-world benchmark designed to evaluate the abilities of LLMs on various rule-guided reasoning tasks. We observe that existing LLMs face significant challenges when they try to tackle problems on RuleArena - even the strongest Claude-3.5 and GPT-4o models can hardly succeed on our hardest tasks. Our further analysis indicates that LLMs struggle to integrate multiple rules or facts cohesively and are prone to irrelevant distractions. RuleArena poses fundamental challenges in complex rule following, and provides a valuable tool for understanding and enhancing the reasoning abilities of LLMs. We envision RuleArena as a foundation for future research to improve LLM performances in solving increasingly complex tasks.
Limitations
While this study provides a comprehensive evaluation and analysis of LLMs’ rule-guided reasoning capabilities, there remain some limitations and numerous promising avenues for future research:
@@
In this work, we only rely on GPT-4o to parse textual responses into structured JSON, facilitating downstream analyses of rule application. A logical next step would be to investigate the use of LLMs for fully automated reasoning evaluations, including the identification of intermediate errors. This direction aligns with the concept of “LLM-as-a-judge” (Zheng et al., 2023), which, despite potential bias or inaccuracies (Xu et al., 2024), offers a scalable alternative to labor-intensive human evaluation and could improve the reliability and granularity of assessment metrics.
@@
Supervised fine-tuning has proven effective in enhancing LLMs’ performance on tasks requiring substantial domain knowledge (Jeong, 2024; Fu et al., 2023; Wu et al., 2024a). While we did not pursue fine-tuning in this study—given the high cost of obtaining extensive rule-guided reasoning data and the limited generalizability to unseen domains—it remains a worthwhile direction. Investigating whether training with datasets from related domains, such as mathematical or logical reasoning tasks or code generation problems, can bolster LLMs’ rule-following performance is an open question worth exploring.
Sophisticated Rule Recall and Aggregation
Our experiments reveal that LLMs frequently struggle with recalling and aggregating the correct rules, and that problem-wise recall strongly correlates with overall accuracy. Addressing these challenges may involve refining rule retrieval mechanisms or integrating structured reasoning frameworks. For instance, approaches that retrieve relevant information dynamically (Trivedi et al., 2023; Zhou et al., 2024) and convert rules into structured data (Pan et al., 2023; Wang et al., 2024) have shown promise in reasoning tasks. Building on these insights, a hybrid system that combines LLM-based reasoning with symbolic reasoners may enhance both the consistency and robustness of rule-guided reasoning in real-world scenarios.